The sum of two numbers is 36. If the larger is divided by the smaller, the quotient is 2 and the remainder is 3. Find the numbers.
Please help.

Answer by mananth(16946)

let the numbers be x & y x>y
x+y = 36

dividend = divisor * quotient + remainder
x= y*2+3
x-2y=3

x+y=36
x-2y=3
multiply first equation by 2 and add to the second
2x+2y=72
x-2y=3
3x=75
/3
x=25
plug value of x in
x+y = 36
25+y=36
y=36-25
y=11
the numbers are 25 & 11